OpenStudy (anonymous):

You have 30 coins. The ratio of heads to tails is 2 to 3. How many coins are heads?

OpenStudy (cwrw238):

2 / (2+3) * 30

OpenStudy (anonymous):

Why? I'm confused with ratios.

OpenStudy (cwrw238):

well there are five parts of the 30 (the ratio is 2: 3 \; 2+3 = 5) heads = 2 and tails = 3 so number heads = 2/5 of 30 tails = 3/5 of 30

OpenStudy (cwrw238):

* (the ratio is 2: 3 so there are 2+ 3 = 5 parts)

OpenStudy (anonymous):

I get it!

OpenStudy (cwrw238):

good

OpenStudy (cwrw238):

so you can now work out number heads right?

OpenStudy (cwrw238):

2/5 * 30

OpenStudy (anonymous):

Yes. # of heads = 12

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

12 is correct

OpenStudy (anonymous):

Ok thanks, seeya